Circular No. 8399

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004ee IN ESO 298-G7 Y.-T. Chen, K. Huang, Z. Lin, H. Lin, and W. Ip, Institute of Astron-

omy, National Central University, Taiwan; and Y. Qiu and J. Hu, National Astronomical Observatories of China, report the discovery of an apparent supernova on unfiltered CCD images taken with the 1-m Lulin Observatory telescope. The new object is located at α = 2^{h}06^{m}14^{s}.62, δ = −37^{o}20′05′′.4 (equinox 2000.0), which is 6′′.0 east and 8′′.0 south of the center of ESO 298-G7 = PGC 8025. Approximate magnitudes for SN 2004ee: 2003 Oct. 30.683 UT, [19.3; 2004 July 31.84, 17.2; Aug. 30.79, 18.1; Sept. 1.834, 18.1.

SUPERNOVA 2004ef IN UGC 12158 T. Boles (cf. IAUC 8397) and M. Armstrong (cf. IAUC 8397) indepen-

dently report their discoveries of an apparent supernova on unfiltered CCD images. Boles’ measured position for the new object is α = 22^{h}42^{m}10^{s}.02, δ = +19^{o}59′40′′.4 (equinox 2000.0), which is ≈ 7′′.2 west and 9′′.2 south of the center of UGC 12158. Approximate magnitudes for SN 2004ef: 2003 Dec. 15 UT, [19.0 (Armstrong); 2004 July 16, [19.5 (Boles); Aug. 19, [19.5 (Boles); Sept. 4.056, 18.5 (Boles); 4.821, 18.5 (Boles); 5.078, 17.5 (Armstrong); 5.861, 17.5 (Armstrong). Position end figures for SN 2004ef by Armstrong are 09^{s}.97, 40′′.1; he adds that nothing is visible at this location on Palomar Sky Survey images from 1990 (limiting red mag 20.8) and 1992 (limiting blue mag 22.5).

SUPERNOVA 2004dj IN NGC 2403 M. K. Argo, T. W. B. Muxlow, R. J. Beswick, and A. Pedlar, Jod-

rell Bank Observatory; and J. M. Marcaide, University of Valencia, report that observations on Aug. 8 with a subset of MERLIN at 4994 MHz give a position for SN 2004dj of α = 7h37m17.044, δ = +65^{o}35′57′′.84 (equinox 2000.0) and a flux density of 1.5 ± 0.15 mJy. The observations were phasereferenced using three calibrators with a resulting positional accuracy of better than 0′′.050. This position is within 0′′.4 of the optical position reported by Itagaki on IAUC 8377 and agrees with the x-ray position reported on IAUC 8390. However, the MERLIN position is 1′′.2 from the VLA position of Stockdale et al. on IAUC 8379, which is possibly affected by inclusion of extended radio emission (IAUC 8391) that is > 3′′.8 from the position of Sandage’s object 96. Continued monitoring of this source is underway.

2004 September 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green